1. A number consists of two digits whose sum is 10 . If 72 be subtracted from the number , the digits are reversed . find the number ? 2. Two kg of sugar and 1/2 kg of tea cost Rs.32 and 3 kg of sugar and 1 kg of tea cost Rs.59. Find the cost of 1 kg of tea 1kg of sugar ?
Answer by checkley71(8403)

1) X+Y=10 & 10X+Y=10Y+X-72 OR 9X-9Y=-72 MULTIPLYING THE FIRST EQUATION BY -9
WE GET
-9X-9Y=-90 ADDING THIS TO
9X-9Y=-72 WE GET
----------
-18Y=-162 OR 18Y=162 OR Y=162/18 OR Y=9 THEREFORE X=10-9 OR X=1
THEREFORE THE ORIGINAL NUMBER IS 91 AND WHEN 72 IS SUBTRACTED WE GET 91-72=19
2) 2S+.5T=.32 & 3S+T=.59 MULTIPLYING THE FIRST EQUATION BY -2 AND ADDING IT TO THE SECOND EQUATION WE GET
-4S-T=-.64
3S+T=.59
---------
-S=-.05 OR S=.05 THEN 2(.05)+.5T=.32 OR .10+.5T=.32 OR .5T=.32-.10 OR .5T=.22
OR T=.22/.5 OR T=44
PROOF 3(.05)+.44=.59 OR .15+.44=.59 OR .59=.59 AND
2(.05)+.5(.44)=.32 OR .10+.22=.32 OR .32=.32